Taco:

Marco can run 3 miles per hour faster than Polo. Marco ran 12 miles in the same time it took Polo to run 8 miles.

Taco:

I need to fill in a table for Marco and Polo's Distance, Rate, and Time. Then determine their speed.

bm717:

@Hero

Hero:

\(\begin{array}{|c|c|c|c|} \hline &\text{Distance}&\text{Rate}&\text{Time}\\ \hline \text{Marco}&&&\\ \hline \text{Polo}&&&\\ \hline \text{Total}&&& \hline \end{array}\)

Taco:

thats what it looks like on my paper

Taco:

umm, so Marco is 12 and Polo is 8

Taco:

Total is 20

Hero:

So did they run the same time or different time?

Taco:

it says for the different distances they ran the same time

Hero:

Correct, they ran the same time so we'll let \(t\) represent the time for both runners: \(\begin{array}{|c|c|c|c|} \hline &\text{Distance}&\text{Rate}&\text{Time}\\ \hline \text{Marco}&12&&t\\ \hline \text{Polo}&8&&t\\ \hline \text{Total}&&& \hline \end{array}\)

Hero:

Okay, so now all we have to do is figure out the rates for both. Who ran faster than whom?

Taco:

Marco

Taco:

uh, 3miles per hour faster than Polo

Hero:

Good, so whatever Polo's speed is, Marco ran 3 miles per hour faster than that:

Hero:

So if we let \(r\) represent Polo's speed, how can we show that Marco ran three miles per hour faster than that? \(\begin{array}{|c|c|c|c|} \hline &\text{Distance}&\text{Rate}&\text{Time}\\ \hline \text{Marco}&12&&t\\ \hline \text{Polo}&8&r&t\\ \hline \text{Total}&&& \hline \end{array}\)

Taco:

r + 3?

Hero:

Correct: \(\color{#0cbb34}{\text{Originally Posted by}}\) @Hero So if we let \(r\) represent Polo's speed, how can we show that Marco ran three miles per hour faster than that? \(\begin{array}{|c|c|c|c|} \hline &\text{Distance}&\text{Rate}&\text{Time}\\ \hline \text{Marco}&12&r+3&t\\ \hline \text{Polo}&8&r&t\\ \hline \text{Total}&&& \hline \end{array}\) \(\color{#0cbb34}{\text{End of Quote}}\)

Hero:

So how can we set up an equation to find \(r\)?

Taco:

r = 8/t ?

Hero:

We need to start off by setting two things that we know are equal, equal to each other.

Hero:

And what would that be @Taco ?

Taco:

the other rate?

Hero:

What two things that we discussed earlier are same for both Marco and Polo?

Taco:

oh time

Taco:

I think I get it

Hero:

Correct, they both ran the same length of time together so we set \(t = t\). Do you think you can set up the next step from here?

Taco:

let me try

Hero:

Post your steps here

Hero:

You can use either the drawing board or the equation buttons located just below the text box.

Taco:

12/(r +3) = 8/r 12r = 8(r +3) 12r = 8r +3 4r = 3 r = 3/4 is that right

Hero:

12r = 8(r +3) Can you try re-doing this step by distributing the 8 using the distributive property? \(a(b + c) = ab + ac\)

Taco:

oh

Taco:

12r = 8(r +3) 12r = 8r + 24 4r = 24 r = 6

Hero:

So what does the 6 represent in this case?

Taco:

Polo runs 6mph ?

Taco:

and Marco runs 9mph?

Hero:

Correct.

Taco:

oh wow

Taco:

thank you I didnt understand this

Hero:

Let's check by making sure the times are the same for both.

Taco:

t = 8/6 = 4/3 t = 12/9 = 4/3

Hero:

Very good. So they both ran about 1 + 1/3 hours. 6 miles or 9 miles in 1 + 1/3 hours. Seems a reasonable time for running mileages.

Taco:

Can you help me with more?

Hero:

Create a new post for the next question.

Taco:

i don't see the text box

Hero:

Because you have to close this question first.

Taco:

oh